OpenStudy (anonymous):

Shureka Washburn has score of 72,67,82, and 79 on her alebra test. Use an inequality to find the scores she must make on the final exam to pass the course with an average of 77 or higher, given that the final exam counts as two test.

OpenStudy (anonymous):

\[\frac{72+67+82+79+2x}{6}\geq 77\] is a start

OpenStudy (anonymous):

\[\frac{300+2x}{6}\geq 77\] \[300+2x\geq 462\] \[2x\geq 162\] \[x\geq 81\]